# Equilibrium of Charges and Differential Equations Solved by Polynomials II

### v1

@@ Abstract
We continue study of equilibrium of two species of 2d coulomb charges (or point vortices in 2d ideal fluid) started in [19]. Although for two species of vortices with circulation ratio $-1$ the relationship between the equilibria and the factorization/Darboux transformation of the Schrodinger operator was established a long ago, the question about similar relationship for the ratio $-2$ remained unanswered. Here we present the answer.
@@ 1 Introduction
In the absence of external field or background flow, equilibrium positions $z_{i}=x_{i}+{\rm i}y_{i}$ of $N$ coulomb charges or point vortices on the plane are roots of the system of $N$ algebraic equations
$\sum_{j=1,j\not=i}^{N}\frac{Q_{i}}{z_{i}-z_{j}}=0,\quad i=1,\dots,N,$ (1)
where $Q_{i}$ is a value of the $i$th charge, or circulation of the $i$th vortex (for introduction to theory of vortex equilibria see e.g. [2, 3, 8, 17, 21, 24, 25] and references therein). Without loss of generality, for a system consisting of two distinct species of $N=l+m$ vortices we can choose
$Q_{i}=\left\{\begin{array}[]{ll}-1,&i=1\dots l\\ \Lambda,&i=l+1\dots l+m\end{array}\right..$
@@
When $\Lambda=1$ (two species with opposite circulations of equal magnitude), eq. (2) becomes
@@
It first appeared and was completely solved in the paper by Burchnall and Chaundy [6]. Later this case emerged in work by Tkachenko [27] in context of vortex equilibrium (“Tkachenko equation”). Bartmann [7] identified (4) with recurrence relation for the Adler-Moser polynomials $P_{n}(z)$
@@
$P_{0}=1,\quad P_{1}=z,\quad P_{2}=z^{3}+s_{1},\quad\dots\quad P_{n}=P_{n}(z,s_{1},\dots,s_{n-1}).$
@@
The rational solutions of the KdV hierarchy
$u_{n}=-2(\log P_{n})^{\prime\prime}$
are potentials of the Schrodinger operators (in what follows $\partial$ stands for differentiation wrt $z$)
@@
In his work [7], Bartmann also considered the case $\Lambda=2$ of (2)
@@
This case was further studied in [19, 9, 10, 26, 17]. Equation (6) can be solved completely [19, 21] using an approach similar to that by Burchnall and Chaundy [6]. Similarly to (5), in the $\Lambda=2$ case, pairs $q_{n},p_{n}$ or $q_{n},p_{n-1}$ solving (6) also satisfy the first-order recurrence relation [7, 19]
@@
However, in difference from the $\Lambda=1$ case, the procedure for constructing solutions by applications of Darboux transformations remained unknown. Demina and Kudryashov observed that solutions of (6) are related to the Sawada-Kotera hierarchy [9]. This led us to the idea to consider the Darboux transformation for differential operators of the third order, since the Lax operator of the Saswada-Kotera hierarchy has this order  [^1] Demina and Kudryahov considered different aspect of the relationship, namely connection with self-similar reductions of Sawada- Kotera and Kaupp-Kupershmidt hierarchies studied earlier by Kudryashov [18] . For more detail on Sawada-Kotera and Kaup-Kupershmidt hierarchies, see e.g. [5, 12, 15, 14] and references therein. . The Darboux transformations for the third-order operators were studied in [5, 23, 4]. To generate solutions of (6) we will introduce “zero level” extension of such transformations.
Before going into detail on the third-order operators, we will briefly recall method of generation of equilibrium configurations through Darboux transformations of the second-order Schrodinger operators (for more detail and review, see e.g. [21]).
@@ 2 Darboux Transformations of Second Order Operators and Equilibrium of Charges
where $A$ and $A^{*}$ are the first order differential operators
@@
Here, function $\varkappa=\varkappa(z)$ is an eigenfunction of $H$ corresponding to the eigenvalue $\eta$
@@
From (9) and (12) it follows that for any $\lambda$
@@
The above identity implies that the transformation
@@
maps eigenfunctions corresponding to eigenvalue $\lambda$ of $H$ to those of $\hat{H}$. Thus, given the “seed” eigenfunction $\varkappa=\psi(z;\eta)$ of $H$, the new operator $\hat{H}$ and its eigenfunctions $\hat{\psi}(z;\lambda)$ can be obtained from $H$ and its eigenfunctions $\psi(x;\lambda)$. The transformation $H\to\hat{H}$, $\psi\to\hat{\psi}$ given by (12,13) is called the Darboux Transformation.
When $\lambda\not=\eta$ the transform (13) of the two dimensional eigenspace of $H$ is also two-dimensional. However, when $\lambda=\eta$, operator $A$ annihilates eigenfunction $\varkappa$ and (13) maps two dimensional eigenspace of $H$ into one-dimentional sub-eigenspace of $\hat{H}$.
Since, according to (10, 12), $A^{*}$ annihilates $1/\varkappa$ and $\hat{H}=AA^{*}+\eta$, we see that $1/\varkappa$ is an eigenfunction of $\hat{H}$ corresponding to the eigenvalue $\eta$. Then, the general solution $\hat{\varkappa}$ of $\hat{H}\hat{\varkappa}=\eta\hat{\varkappa}$ is
@@
where $C$ is an arbitrary constant, and the primitive of $\varkappa^{2}$ includes another arbitrary constant of integration. This extends the Darboux transformation (13) of an eigenfunction to the case $\lambda=\eta$. Transformation (14) (together with (12)) is usually considered for $\eta=0$ and is called the Darboux transformation at zero energy level.
The Adler-Moser polynomials can be generated by iterative applications of (14). Indeed, rewriting the Tkachenko equation (4) in the Schrodinger form
$\left(-\partial^{2}-2(\log q)^{\prime\prime}\right)\left[\frac{p}{q}\right]=0,$
we can apply zero-enegy level Darboux transformation (12, 14) to $\varkappa=p/q$ and $u=-2(\log q)^{\prime\prime}$. In this way we construct the sequence $\varkappa_{n}=P_{n+1}/P_{n}$, $u_{n}=-2(\log P_{n})^{\prime\prime}$ starting from $\varkappa_{0}=z$ and $u_{0}=0$ (free Schrodinger operator $H_{0}=-\partial^{2}$). Indeed, rewriting (14) in differential form with
@@
we get the first-order recurrence relation for the Adler-Moser polynomials (5). Therefore, given the initial conditions $P_{0}=1$, $P_{1}=z$, the Darboux transformation (14,15) generates all possible equilibrium configurations of charges with $Q_{i}\in\{-1,1\}$.
@@ 3 $\Lambda=2$ Case and Third-Order Operators
etc, where $t_{i}$ and $s_{i}$ stand for arbitrary complex constants  [^2] We set constant corresponding to shift of $z$ to zero, so, without lost of generality, $q_{1}=p_{-1}=z$ emerging in course of integrations of Abel identities.
Now we are ready to look for Darboux transformation in $\Lambda=2$ case: Similarly to the $\Lambda=1$ case, the bilinear equation (6), corresponding to $\Lambda=2$ case can be rewritten in the Schrodinger form
@@
However, we cannot apply Darboux transformations to the above second order-operators, because now $u_{n}=-6(\log q_{n})^{\prime\prime}$ (cf. (12)). To proceed, we note that according to the Abel identities (7), $p_{n}$ in (19, 20) can be replaced with $q_{n+1}^{\prime}q_{n}-q_{n+1}q_{n}^{\prime}$ and we get the third-order differential equation
@@ 4 Darboux Transformations for Third-Order Operators
Equation (26) has a solution $\hat{f}=-f$ and the Darboux-type transform $\hat{L}=\hat{B}\hat{A}$ is obtained from $L=BA$ by the involution [^3] This is similar to the Schrodinger operator case where the Darboux transformation is the involution $A=\partial-f\leftrightarrow A^{*}=-(\partial+f)$, cf (10). $f\to-f$. From (23, 24) it follows that for such a transformation
@@
where the primitives include arbitrary constants of integration. Thus, we have found the zero-level Darboux transformations (27, 30) for the third-order operators of the type $\partial^{3}-u\partial$.
@@ 5 $\Lambda=2$ Configurations and Darboux Transformations
Substituting (31,32) into transformation (30), or equivalently, substituting $\hat{\phi}=p/q^{2}$, $(\varkappa^{-1})^{\prime}=-\varkappa^{-2}p_{-}/q_{-}^{2}$ and $\varkappa=q/q_{-}$ into (29) and rewriting result in the differential form, we obtain
@@
Equations (33, 34) are, modulo constant factors, the first-order recurrence relations (7) with $q_{+}=q_{n+1}$, $q=q_{n}$, $q_{-}=q_{n-1}$ and $p=p_{n}$, $p_{-}=p_{n-1}$. Together with initial conditions $q_{0}=p_{0}=1$, they produce solutions of (6). There also exist other initial conditions for which solutions of (7) are polynomials that are not solutions of (6). They will be considered in the next section.
@@ 6 Darboux Transformations and Terminating Configurations
As was mentioned in section 2, the first-order recurrence relation (5) for $P_{n}$ follows from the Darboux transformation (14, 15). Bilinear equation (4) is not involved in this derivation. Therefore, to get a sequence of equilibrium configurations, the initial conditions for (5) need not necessarily to be those for the Adler-Moser polynomials $P_{0}=1$, $P_{1}=z$. A family of sequences with other initial conditions was introduced in work by Duistermmat and Grunbaum [11] in the context of bi-spectral problem. In relation to vortex equilibria this family appeared in work by O’Neil and Cox-Steib [26] (also see [17]). There, sequences start from $P_{0}=1$, $P_{1}=z^{1/2}$. In difference from the Adler-Moser case, these sequences terminate as a non-factorizable $\varkappa$ emerges at some step of recurrence. Examples of first several sequences are:
@@
Similar situation occurs in the case of the third-order operators. As we saw in section 5, the Darboux transformations (30, 35), together with (36), result in the first-order recurrence relations (7). For two species of charges initial conditions are $p_{0}=q_{0}=1$. However, there exist other initial conditions, where terminating sequences of equilibrium configurations, that are not solutions of (6), are generated by the Darboux transformations. Such sequences of solutions were found in [26] (also see [9], [17]). For example, for $q_{0}=1$, $p_{0}=z^{2}$ the first several sequences are

### v2

@@ Abstract
We continue study of equilibrium of two species of 2d coulomb charges (or point vortices in 2d ideal fluid) started in [20]. Although for two species of vortices with circulation ratio $-1$ the relationship between the equilibria and the factorization/Darboux transformation of the Schrodinger operator was established a long ago, the question about similar relationship for the ratio $-2$ remained unanswered. Here we present the answer: One has to consider Darboux-type transformations of third order differential operators rather than second order Schrodinger operators. Furthermore, we show that such transformations can also generate equilibrium configurations where an additional charge of a third specie is present. Relationship with integrable hierarchies is briefly discussed.
@@ 1 Introduction
Systems of point vortices are weak solutions of the Euler equation for a flow of two-dimensional ideal fluid. Steady irrotational flows, which are singular only at a finite number of points are of a particular interest. When such a flow occupies the whole plane and its velocity vanishes at infinity, positions of singularities, i.e. “positions of vortices”, $z_{i}=x_{i}+{\rm i}y_{i}$ are roots of the system of $N$ algebraic equations
$\sum_{j=1,j\not=i}^{N}\frac{Q_{j}}{z_{i}-z_{j}}=0,\quad i=1,\dots,N,$ (1)
where $N$ is the number of vortices and $Q_{i}$ is a value of circulation of the $i$th vortex. Derivation of (1) from the Euler equation can be found e.g. in [22]. For a general introduction to vortex dynamics and equilibrium, see e.g. [2, 3, 9, 18, 22, 25, 26] and references therein.
System (1) has another interpretation: It describes equilibrium of $N$ charges interacting through 2d coulomb potential. Solutions of (1) are stationary points of the coulomb energy function $E=-\sum_{i<j}Q_{i}Q_{j}\log|z_{i}-z_{j}|$. In this interpretation $Q_{i}$ are values of the coulomb charges.
System (1) is invariant under rigid motions and scaling of the plane. From invariance of the stationary value of the coulomb energy under scaling of the plane it follows that the system has solutions only if $\sum_{i<j}Q_{i}Q_{j}=0$. O’Neil showed [25] that for almost all values of $Q_{i}$, satisfying the above condition, there are exactly $(N-2)!$ distinct configurations, i.e. distinct, modulo rigid motions and scaling of the plane, solutions of (1). However, there are “resonant” values of $Q_{i}$, where number of configurations is infinite.
In this paper we mainly deal with two families of resonant cases [^2] For more details on resonant cases see e.g. [26, 29] and references therein. i) two species of vortices with the circulation ratio -1, and ii) two species with the circulation ratio -2. These families are infinite sequences of configurations depending on increasing number of continuous free parameters. They are related to rational solutions of hierarchies of integrable PDEs. Families with an additional vortex of the third specie are considered in Section 6.
To set the problem, we first need to outline history of the question and introduce notations.
Without loss of generality, for a system consisting of two distinct species of $N=l+m$ vortices with circulation ratio $-\Lambda$, we can choose
$Q_{i}=\left\{\begin{array}[]{ll}-1,&i=1\dots l\\ \Lambda,&i=l+1\dots l+m\end{array}\right..$ (2)
@@
In the first resonant case circulation ratio equals -1, $\Lambda=1$ (two species with opposite circulations of equal magnitude, see (2)) and Eq. (3) becomes
@@
It first appeared and was completely solved in the paper by Burchnall and Chaundy [6]. Later this equation emerged in work by Tkachenko [28] in context of vortex equilibrium (“Tkachenko equation”). Bartmann [8] identified bilinear equation (5) with recurrence relation for the Adler-Moser polynomials $P_{n}(z)$
@@
$\begin{array}[]{l}P_{0}=1,\\ P_{1}=z,\\ P_{2}=z^{3}+s_{1},\\ P_{3}=z^{6}+5s_{1}z^{3}+s_{2}z-5s_{1}^{2},\\ \dots\\ P_{n}=P_{n}(z,s_{1},\dots,s_{n-1}),\\ \dots\end{array}$
Examples of several first configurations of this family are shown on Figure 1.
Figure 1: $\Lambda=1$ case: Configurations corresponding to subsequent pairs $(P_{0},P_{1})$, $(P_{1},P_{2})$ and $(P_{2},P_{3})$ of the Adler-Moser polynomials. Roots of the first polynomial in the pair (positions of charges -1) are shown by the gray squares, while roots of the second polynomial in the pair (positions of charges +1) are shown by black circles. In this example $s_{1}=1,s_{2}=1+5{\rm i}$.
@@
The rational solutions of the KdV hierarchy are proportional to the second logarithmic derivative of the Adler-Moser polynomials [1]
$u_{n}=-2(\log P_{n})^{\prime\prime}.$
They are potentials of the Schrodinger operators (in what follows $\partial$ stands for differentiation wrt $z$)
@@
The Darboux transformation method allows to represent solutions of (5) in the Wronskian determinant form [1, 6]. It also allows to construct overcomplete rings of commuting differential operators that contain $H_{n}$ [6, 7].
In his work [8], Bartmann also considered the second resonant case (circulation ratio -2). According to (2), it corresponds to specification $\Lambda=2$ of Eq. (3)
@@
[…]
@@
However, in difference from the $\Lambda=1$ case, the procedure for constructing solutions by applications of Darboux transformations remained unknown.
Krishnamurthy et al. [18] found an integral transformation that produces all the above mentioned resonant families (among other examples). When $\Lambda=1$, it reduces to the Darboux transformation. When $\Lambda=2$, it is a composition of the Darboux transformation and exponentiation (see [22] for details), but not a Darboux-type transformation for a differential equation.
Demina and Kudryashov observed that solutions of (7) are related to the Sawada-Kotera hierarchy [10]. This led us to the idea to consider the Darboux transformation for differential operators of the third order, since the Lax operator of the Sawada-Kotera hierarchy has this order  [^3] Demina and Kudryahov considered different aspect of the relationship, namely connection with self-similar reductions of Sawada- Kotera and Kaup-Kupershmidt hierarchies studied earlier by Kudryashov [19]. . The Darboux transformations for the third-order operators were studied in [4, 5, 24]. To generate solutions of (7), we will introduce “zero level” extension of such transformations. Additionally, we will show that apart from solutions of (7), configurations with an additional charge of the third specie (previously found by O’Neil and Cox-Steib in [27]) can be also constructed with help of introduced transformations.
Despite similarities between the $\Lambda=1$ and $\Lambda=2$ cases, the fact that the $\Lambda=2$ case is related to Darboux transformations of differential operators of the third order, rather than those of the second order Schrodinger operators, implies some qualitative difference with the $\Lambda=1$ case from the vortex equilibrium viewpoint. In particular, translating configurations, which are solutions of a generalization of (1), cannot be now constructed by the Darboux transformation method.
Another unlikeness between the cases arises when one looks for determinant representations of configurations. These and other dissimilarities are discussed in the conclusion section. There, we also briefly go into relationships between the $\Lambda=2$ configurations and rational solutions of the Sawada-Kotera and Kaup-Kupershmidt hierarchies in context of Pfaffian/determinant representations of these configurations.
Before going into detail on the third-order operators, we will recall method of generation of equilibrium configurations through Darboux transformations of the second-order Schrodinger operators (for more detail and review, see e.g. [22]).
@@ 2 Darboux Transformations of Second Order Operators and Equilibrium of Charges
where $A$ and $A^{*}$ are (formally conjugated) first order differential operators
@@
Here, the function $\varkappa=\varkappa(z)$ is an eigenfunction of $H$ corresponding to eigenvalue $\eta$
@@
From (10) and (13) it follows that for any constant $\lambda$
@@
Acting by both sides of the above identity on the eigenfunction $\psi=\psi(z;\lambda)$ of $H$: $H\psi=\lambda\psi$, we come to the conclusion that the transformation
@@
maps eigenfunctions corresponding to eigenvalue $\lambda$ of $H$ to those of $\hat{H}$. Thus, given the seed eigenfunction $\psi(z;\eta)=\varkappa$ corresponding to the seed eigenvalue $\eta$ of $H$ (see (12)), the new operator $\hat{H}$ and its eigenfunctions $\{\hat{\psi}=\hat{\psi}(z;\lambda)$ : $\hat{H}\hat{\psi}=\lambda\hat{\psi}\}$ can be obtained from $H$ and its eigenfunctions $\{\psi=\psi(z;\lambda)$ : $H\psi=\lambda\psi\}$. The transformation $H\to\hat{H}$, $\psi\to\hat{\psi}$ given by (13, 14) is called the Darboux transformation.
For any eigenvalue, except the seed eigenvalue $\eta$, i.e. for $\lambda\not=\eta$, the transform (14) of the two dimensional eigenspace of $H$ is also two-dimensional. However, when $\lambda=\eta$, operator $A$ annihilates the seed eigenfunction $\psi(z;\eta)=\varkappa$ and (14) maps two dimensional eigenspace of $H$ into one-dimentional sub-eigenspace of $\hat{H}$.
Since, according to (11), $A^{*}$ annihilates $1/\varkappa$, and, by (13), $\hat{H}=AA^{*}+\eta$, we see that $1/\varkappa$ is an eigenfunction of $\hat{H}$ corresponding to the eigenvalue $\eta$, i.e $\hat{H}[1/\varkappa]=\eta/\varkappa$. Then, the general solution $\hat{\varkappa}$ of the second-order differential equation $\hat{H}\hat{\varkappa}=\eta\hat{\varkappa}$ can be obtained from its particular solution $1/\varkappa$ by elementary methods. It equals
@@
where $C$ is an arbitrary constant, and the primitive of $\varkappa^{2}$ includes another arbitrary constant of integration. This extends the Darboux transformation (14) of an eigenfunction to the case $\lambda=\eta$. Transformation (15) is usually considered for $\eta=0$ and is called the Darboux transformation at zero energy level.
Darboux transformations can be iterated. Resulting transforms depend on seed eigenvalues $\eta=\eta_{1},\eta=\eta_{2},\dots$ chosen at different iterations. The Adler-Moser polynomials can be generated by such iterations, where the same seed eigenvalue $\eta_{i}=0$ is taken at each step. Indeed, rewriting the Tkachenko equation (5) in the Schrodinger form
$\left(-\partial^{2}-2(\log q)^{\prime\prime}\right)\left[\frac{p}{q}\right]=0,$ (16)
we can apply zero-enegy level Darboux transformation (13, 15) to $\varkappa=p/q$ and $u=-2(\log q)^{\prime\prime}$ (see Eqs. (9, 12) for $\eta=0$). In this way we construct the sequence $\varkappa_{n}=P_{n+1}/P_{n}$, $u_{n}=-2(\log P_{n})^{\prime\prime}$ starting from $\varkappa_{0}=z$ and $u_{0}=0$ (free Schrodinger operator $H_{0}=-\partial^{2}$). Indeed, rewriting (15) in differential form with
@@
we get the first-order recurrence relation for the Adler-Moser polynomials (6). Therefore, given the initial conditions $P_{0}=1$, $P_{1}=z$, the Darboux transformation (15,17) generates all possible equilibrium configurations of charges with $Q_{i}\in\{-1,1\}$.
@@ 3 $\Lambda=2$ Case and Third-Order Operators
etc, where $t_{i}$ and $s_{i}$ stand for arbitrary complex constants  [^4] We set constant corresponding to shift of $z$ to zero, so, without loss of generality, $q_{1}=p_{-1}=z$. emerging in course of integrations of Abel identities. Examples of several first equilibrium configurations of the $n\geq 0$ sequence are shown on Figure 2.
Figure 2: $\Lambda=2$ case: Configurations corresponding to subsequent pairs $(q_{1},p_{0})$, $(q_{1},p_{1})$ and $(q_{2},p_{1})$. Roots of the first polynomial in the pair (positions of charges +2) are shown by the black circles, while roots of the second polynomial in the pair (positions of charges -1) are shown by gray squares. In this example $t_{1}=1,s_{2}=1+5{\rm i}$.
Now we are ready to look for Darboux transformation in $\Lambda=2$ case: Similarly to the $\Lambda=1$ case, the bilinear equation (7), corresponding to $\Lambda=2$ case, can be rewritten in the Schrodinger form
@@
However, we cannot apply Darboux transformations to the above second order operators, because now $u_{n}=-6(\log q_{n})^{\prime\prime}$ (and not $-2(\log q_{n})^{\prime\prime}$, cf. (13,16)). To proceed, we note that according to the Abel identities (8), $p_{n}$ in (21, 22) can be replaced with $q_{n+1}^{\prime}q_{n}-q_{n+1}q_{n}^{\prime}$ and we get the third-order differential equation
@@ 4 Darboux Transformations for Third-Order Operators
Equation (28) has a solution $\hat{f}=-f$ and the Darboux-type transform $\hat{L}=\hat{B}\hat{A}$ is obtained from $L=BA$ by the involution [^5] This is similar to the Schrodinger operator case where the Darboux transformation is the involution $A=\partial-f\leftrightarrow A^{*}=-(\partial+f)$, cf (11). $f\to-f$. From (25, 26) it follows that for such a transformation
@@
where the primitives include arbitrary constants of integration. Thus, we have found the zero-level Darboux transformations (29, 32) for the third-order operators of the type $\partial^{3}-u\partial$.
@@ 5 $\Lambda=2$ Configurations and Darboux Transformations
Substituting (33,34) into transformation (32), or equivalently, substituting $\hat{\phi}=p/q^{2}$, $(\varkappa^{-1})^{\prime}=-\varkappa^{-2}p_{-}/q_{-}^{2}$ and $\varkappa=q/q_{-}$ into (31) and rewriting result in the differential form, we obtain
@@
Equations (35, 36) are, modulo constant factors, the first-order recurrence relations (8) with $q_{+}=q_{n+1}$, $q=q_{n}$, $q_{-}=q_{n-1}$ and $p=p_{n}$, $p_{-}=p_{n-1}$. Together with initial conditions $q_{0}=p_{0}=1$, they produce solutions of (7). There also exist other initial conditions for which solutions of (8) are polynomials that are not solutions of (7). They will be considered in the next section.
@@ 6 Darboux Transformations and Terminating Configurations
As was mentioned in section 2, the first-order recurrence relation (6) for $P_{n}$ follows from the Darboux transformation (15, 17). Bilinear equation (5) is not involved in this derivation. Therefore, to get a sequence of equilibrium configurations, the initial conditions for (6) need not necessarily to be those for the Adler-Moser polynomials $P_{0}=1$, $P_{1}=z$. A family of sequences with other initial conditions was introduced in work by Duistermmat and Grunbaum [12] in the context of bi-spectral problem. In relation to vortex equilibria this family appeared in work by O’Neil and Cox-Steib [27] (also see [18]). There, sequences start from $P_{0}=1$, $P_{1}=z^{1/2}$. In difference from the Adler-Moser case, these sequences terminate as a non-factorizable $\varkappa$ emerges at some step of recurrence. Examples of first several sequences are:
@@
Similar situation occurs in the case of the third-order operators. As we saw in section 5, the Darboux transformations (32, 37), together with (38), result in the first-order recurrence relations (8). For two species of charges initial conditions are $p_{0}=q_{0}=1$. However, there exist other initial conditions, where terminating sequences of equilibrium configurations, that are not solutions of (7), are generated by the Darboux transformations. Such sequences of solutions were found in [27] (also see [10], [18]). For example, for $q_{0}=1$, $p_{0}=z^{2}$ the first several sequences are